OpenStudy (dallaslifebaby01):

which of the following pairs of ratios forms a proportion? 1) 8/9, 12/18 2) 2.3/3.4, 3.0/3.6 3) 21.1/ 14.4, 1.1/1.2 4) 5/2, 4/1.6

OpenStudy (princeharryyy):

proportion => a/b =b/c

OpenStudy (princeharryyy):

It can be d. but directly no one.

OpenStudy (dallaslifebaby01):

okay thank you!

OpenStudy (princeharryyy):

:)

jhonyy9 (jhonyy9):

the last one than you simplify the second fraction will be 4/1,6 = 2/0,8 so 5/2 and 2/0,8 form a proportion in this way hope helped